# A Formal Descriptive Language for Learning Dynamics A Five-Layer Structural Coordinate System

### v1

@@ 2.1 Cognitive Load Theory and Instructional Design
Cognitive Load Theory (CLT) has provided a highly influential set of principles for instructional design by foregrounding the limitations of working memory and by distinguishing among intrinsic, extraneous, and germane cognitive load (Sweller1988CognitiveLoad; Sweller2019CognitiveArchitecture). Within CLT, these distinctions function primarily as design-relevant categories, offering guidance on how instructional materials and task representations should be structured in order to reduce unnecessary processing demands and to support learning-relevant activity. Importantly, later formulations of the theory emphasize that intrinsic load is not a fixed property of instructional materials themselves, but depends on the interaction between task structure and the learner’s prior knowledge, while extraneous and germane load are defined in terms of how information is presented and processed rather than as inherent attributes of stimuli.
@@ 2.2 Guidance, Structure, and Failure as a Learning Condition
Debates on instructional guidance and the literature on productive failure jointly illustrate that learning cannot be characterized as a monotone function of ease, guidance, or observable success. The critique of minimal guidance emphasizes that unguided or weakly guided environments can impose excessive cognitive load on novice learners, often consuming limited working-memory resources without yielding durable learning (Kirschner2006MinimalGuidance). In contrast, the productive failure line of work has demonstrated that initial struggle or failure can, under certain conditions, contribute positively to deeper learning outcomes (Kapur2016ProductiveFailure). Subsequent studies have further clarified that such benefits are highly conditional, depending on factors such as task structure, opportunities for consolidation, and the learner’s interaction with the environment (SinhaKapur2019PF).
@@ 2.3 Motivation and Learner Experience
Motivational theories such as Self-Determination Theory (SDT) provide a rich account of autonomy, competence, and relatedness as determinants of sustained engagement (DeciRyan2000SDT). These constructs capture essential aspects of learner experience, but they are often introduced as explanatory variables at a different descriptive level than the time evolution of understanding or capability. As a result, motivational states can become “global labels” that obscure where, structurally, a learning process becomes self-reinforcing or collapses.
@@ 2.4 Dynamical Systems Approaches to Cognition and Development
Dynamical systems perspectives have long argued that cognition and development should be understood as time-evolving processes characterized by non-linearity, multi-timescale interactions, and emergent stability (ThelenSmith1994DynamicSystems; vanGelder1998DynamicalHypothesis). These approaches are conceptually aligned with the present work in treating learning as dynamics rather than as a static state.
@@ 2.5 Learning Sciences as a Reference Frame and Positioning
Contributions from the learning sciences provide an integrative reference frame that spans cognition, instruction, and context, emphasizing that learning outcomes arise from interactions among prior knowledge, activity, environment, and social organization (Bransford2000HowPeopleLearn). A central message of this tradition is that learning with understanding cannot be reduced to properties of tasks or instructional materials alone, but must be analyzed in relation to the learner’s existing knowledge structures and ongoing activity.
@@ 7 Future Work
Most notably, the internal structure of the subjective evaluation interface $R$ is treated here as an abstract regulatory interface responding to learning dynamics, without commitment to specific reward semantics or motivational constructs. Empirical and theoretical considerations suggest that such evaluation may admit multiple control modes with distinct temporal and functional characteristics—such as sensitivity to change, sensitivity to state, and attenuation under sustained uncertainty. Articulating these modes remains a task for future work, to be pursued in a manner that preserves the descriptive and non-normative character of the framework.

### v2

@@ 2.1 Cognitive Load Theory and Instructional Design
Cognitive Load Theory (CLT) has provided a highly influential set of principles for instructional design by foregrounding the limitations of working memory and by distinguishing among intrinsic, extraneous, and germane cognitive load (Sweller, 1988; Sweller et al., 2019). Within CLT, these distinctions function primarily as design-relevant categories, offering guidance on how instructional materials and task representations should be structured in order to reduce unnecessary processing demands and to support learning-relevant activity. Importantly, later formulations of the theory emphasize that intrinsic load is not a fixed property of instructional materials themselves, but depends on the interaction between task structure and the learner’s prior knowledge, while extraneous and germane load are defined in terms of how information is presented and processed rather than as inherent attributes of stimuli.
@@ 2.2 Guidance, Structure, and Failure as a Learning Condition
Debates on instructional guidance and the literature on productive failure jointly illustrate that learning cannot be characterized as a monotone function of ease, guidance, or observable success. The critique of minimal guidance emphasizes that unguided or weakly guided environments can impose excessive cognitive load on novice learners, often consuming limited working-memory resources without yielding durable learning (Kirschner et al., 2006). In contrast, the productive failure line of work has demonstrated that initial struggle or failure can, under certain conditions, contribute positively to deeper learning outcomes (Kapur, 2016). Subsequent studies have further clarified that such benefits are highly conditional, depending on factors such as task structure, opportunities for consolidation, and the learner’s interaction with the environment (Sinha and Kapur, 2019).
@@ 2.3 Motivation and Learner Experience
Motivational theories such as Self-Determination Theory (SDT) provide a rich account of autonomy, competence, and relatedness as determinants of sustained engagement (Deci and Ryan, 2000). These constructs capture essential aspects of learner experience, but they are often introduced as explanatory variables at a different descriptive level than the time evolution of understanding or capability. As a result, motivational states can become “global labels” that obscure where, structurally, a learning process becomes self-reinforcing or collapses.
@@ 2.4 Dynamical Systems Approaches to Cognition and Development
Dynamical systems perspectives have long argued that cognition and development should be understood as time-evolving processes characterized by non-linearity, multi-timescale interactions, and emergent stability (Thelen and Smith, 1994; van Gelder, 1998). These approaches are conceptually aligned with the present work in treating learning as dynamics rather than as a static state.
@@ 2.5 Learning Sciences as a Reference Frame and Positioning
Contributions from the learning sciences provide an integrative reference frame that spans cognition, instruction, and context, emphasizing that learning outcomes arise from interactions among prior knowledge, activity, environment, and social organization (Bransford et al., 2000). A central message of this tradition is that learning with understanding cannot be reduced to properties of tasks or instructional materials alone, but must be analyzed in relation to the learner’s existing knowledge structures and ongoing activity.
@@ 3.5 Layer 4: Subjective Evaluation Interface
The structural necessity of this design choice is examined in Appendix A, where it is shown that alternative couplings—most notably strictly state-dependent evaluation—generically induce collapse of exploration and loss of sustained learning through the geometry of the evaluation–exploration loop itself. The present formulation $r=R\!\left(\frac{dx}{dt},E\right)$ thus arises not from normative assumptions about learning objectives, but from structural constraints on the stability of adaptive dynamics.
@@ 7 Future Work
Most notably, the internal structure of the subjective evaluation interface $R$ is treated here as an abstract regulatory interface responding to learning dynamics, without commitment to specific reward semantics or motivational constructs. The analysis in Appendix A establishes wiring-level constraints on evaluation; a distinct open question is how the internal structure of $R$ realizes multiple modes and timescales while respecting those constraints. Empirical and theoretical considerations suggest that such evaluation may admit multiple control modes with distinct temporal and functional characteristics—such as sensitivity to change, sensitivity to accumulated change, and attenuation under sustained uncertainty. Articulating these modes remains a task for future work, to be pursued in a manner that preserves the descriptive and non-normative character of the framework.
@@
Appendix A Structural Stability Constraints on Evaluation and Representation Selection
In this appendix, we analyze how differences in the coupling among evaluation, exploration, and internal dynamics affect the stability of learning, thereby clarifying the structural constraints underlying the evaluation interface introduced in Section 3.5.
Here, we consider an adaptive learning system in which the continuous change of the internal state $x$, $\frac{dx}{dt}$, and the interaction with the environment $E$ are controlled by an input acquisition policy. The evaluation $R$ is, in principle, a two-argument evaluation interface that depends on change in the internal state and feedback from the environment, thereby determining the evaluation signal $r$:
$r=R\!\left(\frac{dx}{dt},E\right).$
In this setting, policy update and input acquisition may occur as discrete acquisition events indexed by $k\rightarrow k+1$:
$r_{k}\xrightarrow{U}\pi_{k+1}\xrightarrow{S}e_{k+1}.$
Here, $\pi$ is the input acquisition policy. $U$ denotes the policy update ($r_{k}\mapsto\pi_{k+1}$). $S$ denotes the input acquisition (sampling) operator ($\pi_{k+1}\mapsto e_{k+1}$).
The internal dynamics $\frac{dx}{dt}$ evolve continuously, and discrete acquisition events function as triggers that shape the subsequent continuous change. However, in order to visualize the events occurring in each loop succinctly, in this appendix we schematically approximate changes in the internal dynamics as discrete. Accordingly, we describe the change in $\frac{dx}{dt}$ mediated across steps by the internally driven loop as $\dot{x}_{k}\rightarrow\dot{x}_{k+1}$. Here, $\dot{x}_{k}$ is used as a representative symbol for $\frac{dx}{dt}$ within each step.
In what follows, we describe the same event by decomposing it into two schematics: a projection onto an internally driven loop and a projection onto a loop that includes interaction with the environment. The internally driven loop proceeds as follows:
$\dot{x}_{k}\xrightarrow{R}r_{k}\xrightarrow{U}\pi_{k+1}\xrightarrow{S}e_{k+1}\xrightarrow{\Phi}(\ell_{x,k+1},\ell_{n,k+1})\xrightarrow{G}\dot{x}_{k+1}.$
In addition, environmental feedback is generated through externalization and enters evaluation:
$x_{k}\xrightarrow{Q}y_{k}\xrightarrow{H}E_{k}\xrightarrow{R}r_{k}\xrightarrow{U}\pi_{k+1}\xrightarrow{S}e_{k+1}\xrightarrow{\Phi}(\ell_{x,k+1},\ell_{n,k+1})\xrightarrow{G}\dot{x}_{k+1}\xrightarrow{\mathcal{I}}x_{k+1}.$
$\Phi$ is a fixed family of representation decomposition mappings. $Q$ denotes externalization ($x_{k}\mapsto y_{k}$). $H$ denotes the environmental response mapping ($y_{k}\mapsto E_{k}$). $G$ is a mapping that determines the internal dynamics (change in the internal state). $\mathcal{I}$ denotes the step-wise state update induced by the continuous-time flow (integration over the step).
In this sense, a single event may admit simultaneous projections of its components as $E$ and $e$ onto multiple functional bases (e.g., evaluation $R$ and acquisition $S$), without implying any serial transformation between them.
A.1 Description as Probabilistic Sampling of $\Phi$-Selection
In this appendix, we describe the selection of $\Phi$ as a process in which the policy $\pi$ acquires an input $e$, a bias over candidate $\Phi$ arises through the response to that $e$, and which $\Phi$ is chosen is then determined probabilistically according to the resulting weights. The probabilistic character here does not assume that an explicit probability distribution is represented inside the learner; rather, it is a descriptive summary of tendencies observed through repeated interaction. Furthermore, as long as the same $e$ is maintained for some period of time, $\Phi$-selection repeatedly occurs for that same $e$: in some cases the same $\Phi$ may continue to be selected, while in other cases different $\Phi$ may be selected. Accordingly, the gentler (less sharply peaked) the bias is, the more diverse $\Phi$ can be selected, making it possible to interpret a single $e$ from multiple perspectives; conversely, the sharper (more peaked) the bias is, the more likely the same $\Phi$ will continue to be selected, and deepening of understanding within the same context will occur. Moreover, this bias is not fixed: it may change over the long term through the accumulation of interactions with $e$ (in particular, through repeated contact with $e$ derived from others’ $\Phi$). However, in this paper we treat such long-term changes in bias not as “learning” but as plastic changes in $\Phi$-selection (“noticing”), and we distinguish them from deepening of the internal state $x$ (“understanding”).
A.2 Structural Degeneracy of State-Dependent Evaluation
In this framework, we do not assume that evaluation takes the internal state $x$ as a direct input. This is because the internal state $x$ has no intrinsic reference scale by which it can be evaluated, even from the learner’s own perspective. Without a standard of comparison, the absolute magnitude of $x$ cannot be internally monitored, whereas the internal change $\dot{x}$ may be experienced as variation over time. Moreover, coupling evaluation directly to $x$ risks conflating state description with control objectives. To make this structural consequence explicit, we temporarily consider the counterfactual wiring $r=R(x)$ and examine the resulting loop geometry. In discrete acquisition events, this is instantiated as $r_{k}=R(x_{k})$.
A.2.1 State-Dependent Evaluation (Strictly Internal Form)
We consider a strictly internal configuration in which evaluation is defined solely as a function of the internal state:
$r=R(x).$
Since the internal state $x$ is defined only relative to a chosen representational decomposition $\Phi$ and has no privileged intrinsic reference scale, state-based evaluation is not invariant under representational change. Consequently, the control objective induced by $R(x)$ is structurally tied not to any external criterion of success but to the maintenance of internal coherence under the currently active representation.
Under this wiring, the learning loop closes through evaluation, policy update, and acquisition, feeding back to internal variation.
A.2.2 Structural Assumptions Underlying Degeneracy of State-Dependent Evaluation
The degeneracy of learning under strictly state-dependent evaluation arises not from any particular optimization objective but from the geometry of coupling among evaluation, policy update, and internal dynamics. The following mild structural assumptions suffice.
Assumption 1 (Evaluation-Preserving Policy Update).
The policy update operator $U$ biases the input acquisition policy $\pi$ in the direction that makes it more likely for interaction patterns under which the evaluation signal $r$ remains high to occur again.
Assumption 2 (Structural Bias of State-Based Evaluation).
When evaluation is defined as $r=R(x)$, the absence of an external reference structurally favors high values of $r$ in regions where internal representations are maintained as coherent and predictive under the currently active decomposition $\Phi$. Unless additional structure is artificially imposed on $R$, state-based evaluation therefore tends to assign persistently high $r$ in the direction that keeps change in the internal state small. This degeneracy can be avoided only by imposing additional structure that defines an evaluation scale invariant under representational changes (which is precisely what the present framework avoids assuming). Accordingly, the collapse described here is not tied to a particular objective, but follows generically from the coupling geometry when evaluation depends strictly on $x$.
Assumption 3 (Controllability of Internal Variation).
Through the composition of the acquisition operator $S$, the representation family $\Phi$, and the internal dynamics $G$, the policy $\pi$ can adjust how the internal state changes by appropriately selecting the input $e$.
Under these conditions, the closed loop
$x_{k}\xrightarrow{R}r_{k}\xrightarrow{U}\pi_{k+1}\xrightarrow{S}e_{k+1}\xrightarrow{\Phi}(\ell_{x,k+1},\ell_{n,k+1})\xrightarrow{G}\dot{x}_{k+1}$
has a structurally stable attractor in the direction that keeps change in the internal state small.
The input acquisition policy $\pi$ converges toward input patterns that maintain persistently high values of the evaluation signal $r$ by suppressing change in the internal state. As a result, a stable behavior emerges in which change in the internal state is suppressed, but learning is degenerate: internal representations remain coherent under the current decomposition $\Phi$, while the system loses the capacity for sustained adaptation.
A.3 Transition-Dependent Evaluation
To avoid the structural collapse induced by state-dependent evaluation, the evaluation interface must be coupled to quantities that are robust to representational change and that reflect actual interaction with the environment. This requirement is satisfied when evaluation is defined as:
$r=R\!\left(\frac{dx}{dt},E\right).$
In discrete acquisition events, we write this schematically as $r_{k}=R(\dot{x}_{k},E_{k})$.
Under this coupling, high values of $r$ cannot be maintained merely by suppressing internal change. Internal state $x$ becomes evaluable only insofar as it is consistent with external consequences $E$, and conversely, environmental feedback becomes meaningful only relative to the internal transformations that produced it.
As a result, the learning loop acquires a fundamentally different geometry:
$x_{k}\xrightarrow{Q}y_{k}\xrightarrow{H}E_{k}\xrightarrow{R}r_{k}\xrightarrow{U}\pi_{k+1}\xrightarrow{S}e_{k+1}\xrightarrow{\Phi}(\ell_{x,k+1},\ell_{n,k+1})\xrightarrow{G}\dot{x}_{k+1}\xrightarrow{\mathcal{I}}x_{k+1}.$
In this configuration, because evaluation constrains internal dynamics through externally grounded consequences, the policy $\pi$ cannot preserve high values of $r$ by suppressing change in the internal state. Therefore, exploration through $S$ persists, the effective distribution over representation decompositions $\Phi$ maintains diversity, and long-term adaptation becomes structurally stable.
When environmental feedback $E$ becomes persistently adverse, the update $U$ may favor acquisition policies that prioritize immediate reduction of adverse consequences, yielding a stabilization regime (a short-term protective mode) in which inputs are sampled to stabilize the $E$-channel rather than to diversify $\Phi$-selection. In such a stabilization regime, sampling tends to concentrate on a narrower subset of interaction patterns (even when the nominal input class $e$ is unchanged), which can sharpen the effective bias over $\Phi$ and make alternative decompositions less likely. Note that long-term sharpening of the effective bias in $\Phi$-selection may arise either from plastic changes in the evaluation interface $R$ (e.g., shifts in sensitivity or gain with respect to $(\dot{x},E)$) or from changes in the effective environmental feedback channel (e.g., via $x\xrightarrow{Q}y\xrightarrow{H}E$); the present framework is compatible with both interpretations.
This structural distinction is independent of any particular optimization objective or learning algorithm. It is a consequence of the coupling geometry among evaluation, exploration, and internal dynamics.
A.4 Conclusion
The stability of learning in adaptive systems is determined not by the form of optimization objectives or the details of learning algorithms but by the coupling geometry among evaluation, exploration, and internal dynamics.
When evaluation is coupled directly to the internal state, $r=R(x)$, the learning loop generically acquires a structurally stable attractor in which internal variation is suppressed, exploration collapses, and sustained learning attenuates. This degeneration arises not from a particular design choice but from the representational dependence of state itself.
By contrast, when evaluation is defined over internal transition and environmental feedback, $r=R\!\left(\frac{dx}{dt},E\right)$, the learning loop is structurally constrained so as to preserve exploration, maintain representational diversity, and support long-term adaptation. In this configuration, learning stability emerges as a geometric property of the evaluation–exploration coupling.
Thus, the persistence of learning is governed not by what is optimized but by how the system is wired. In this sense, this framework characterizes learning not as objective maximization but as a problem of structural coupling. The degeneracy result holds for any learning architecture satisfying Assumptions 1–3, independent of parameterization or learning rule.